# Smoother-type a posteriori error estimates for finite element methods

### v2

@@ Abstract
This work develops user-friendly a posteriori error estimates of finite element methods, based on smoothers of linear iterative solvers. The proposed method employs simple smoothers, such as Jacobi or Gauss–Seidel iteration, on an auxiliary finer mesh to process the finite element residual for a posteriori error control. The implementation has linear complexity and requires only a coarse-to-fine prolongation operator. For symmetric problems, we prove the reliability and efficiency of smoother-type error estimators under a saturation assumption. Numerical experiments for various PDEs demonstrate that the proposed smoother-type error estimators outperform residual-type estimators in accuracy and exhibit robustness with respect to parameters and polynomial degrees.
@@ 1 Introduction
A posteriori error estimation of Finite Element Methods (FEMs) is also closely related to iterative linear solvers because both methods aim at developing computable approximations to the action of solution operators on some residuals. For instance, the idea of hierarchical basis solvers motivated the development of hierarchical basis error estimators in [9, 6]. Multigrid solvers combined with superconvergence are used for constructing a posteriori error estimates in [12, 13]. It has been shown in [51, 52] that additive operator preconditioning on infinite-dimensional spaces yields implicit and parameter-robust a posteriori error estimates for $H({\rm grad})$, $H({\rm curl})$ and $H({\rm div})$ problems. Conversely, a posteriori error estimation has been applied to design coarse spaces in multigrid methods and to control iterative algebraic errors in domain decomposition methods, see [68, 44, 15].
@@
We shall follow the research line in [51, 52] to derive a posteriori error estimates based on smoothers of iterative linear solvers. Smoothers are simple basic iterations such as Jacobi or Gauss–Seidel method on a fine mesh and are efficient for damping the high frequency component of the FEM error. We shall show that appropriate smoothers directly yield high-quality a posteriori error estimates of FEMs for various PDEs such as the Poisson, curl-curl, Biot, convection-diffusion and Helmholtz equations. In numerical experiments, smoother-based error estimators exhibit sharper effectivity ratio than residual-type a posteriori error estimates. In contrast to the additive preconditioning-type implicit error estimators derived in [51, 52], which involve incomputable local problems, the proposed smoother-type a posteriori error estimators are fully computable. Moreover, this work introduces a new family of error estimators inspired by contractive iteration methods, which are particularly well-suited for handling non-symmetric problems.
@@
The implementation of the smoother-type a posteriori error estimates on a mesh $\mathcal{T}_{h}$ is similar to a two-grid solver. In particular, we shall make use of a simple smoother $S$, e.g., a Jacobi, block Jacobi or Gauss–Seidel iteration on an auxiliary finer mesh $\mathcal{T}_{h/2}$. Then a smoother-type a posteriori error estimate is of the form $\|Sr\|$ or $\sqrt{\langle r,Sr\rangle}$, where $r$ is the finite element residual transferred to $\mathcal{T}_{h/2}$, $\|\bullet\|$ is the norm of interest and $\langle\bullet,\bullet\rangle$ the pairing between the primal and dual spaces. The reliability and efficiency of our a posteriori error estimates are confirmed under a saturation assumption. Using the same assumption, adaptive finite and boundary element methods driven by the difference $\|u_{h}-u_{h/2}\|$ of finite element solution on nested meshes have been analyzed in [39, 38], while computing finer mesh solution $u_{h/2}$ is much more expensive than the smoother action $Sr$.
@@
The rest of the paper is organized as follows. In Section 2 we provide preliminaries and set up the abstract framework. In Section 3 we employ smoothers of iterative solvers to derive a posteriori error estimates for positive-definite problems. Section 4 is devoted to the smoother-type a posteriori error estimates of symmetric saddle-point systems. Section 5 illustrates the applications of smoother-type estimators to convection-diffusion and Helmholtz equations. Section 6 is the concluding remark.
@@ 2.1 Abstract Error Estimators
Throughout this paper, $V_{h}$ is a proper finite element space based on a partition $\mathcal{T}_{h}$ of the underlying domain $\Omega$ for (2.1a). The form $a$ includes a bounded linear operator $A:V\rightarrow V^{*}$ defined as follows:
@@
We adopt the notation $C_{1}\lesssim C_{2}$ (resp. $C_{1}$ $\lesssim C_{2}$) provided $C_{1}\leq C_{3}C_{2}$ with $C_{3}$ being a generic constant independent of quantities of interest, e.g., the mesh size parameter $h$ and target functions. We say $C_{1}\eqsim C_{2}$ provided $C_{1}\lesssim C_{2}$, $C_{2}\lesssim C_{1}$. This notation generalizes to comparison between SPD operators in an obvious way. It has been observed in [51] that additive two-level preconditioning yields a reliable and efficient error estimator for SPD problems. For completeness, we include a brief proof.
@@
We remark that $\langle r,S_{a}r\rangle^{\frac{1}{2}}$ is applicable to SPD problems as well as symmetric but indefinite systems posed in a Cartesian product space $V_{1}\times V_{2}\times\cdots$, where $S_{a}={\rm diag}(S_{V_{1}},S_{V_{2}},\ldots)$ is block diagonal with each $S_{V_{i}}$ being a SPD smoother on $V_{i}$, see Section 4. However, $\langle r,S_{a}r\rangle^{\frac{1}{2}}$ could not be applied to strongly non-symmetric problems because $\langle r,S_{a}r\rangle$ is not necessarily nonnegative in such cases. Our new observation is that a contractive iterative solver for $Au=f$
@@
When AA is the SPD operator induced by the inner product of $H^{1}(\Omega)$, $H({\rm curl},\Omega)$, $H({\rm div},\Omega)$ or their finite element variants, the iteration (2.7) with $S_{m}$ in Algorithm 2.1 given by a stable decomposition is uniformly contractive under the norm $\|\bullet\|_{A}$, i.e., $\|I-B_{m}A\|_{A}<1$ is independent of the mesh size parameter $h$.
@@ 2.2 Application to the Poisson Equation
where $V_{h}\subset H_{0}^{1}(\Omega)$ is a finite element space of continuous and piecewise $\mathcal{P}_{1}$ polynomials. Let $\{a_{h,k}\}_{1\leq k\leq N_{h}}$ be the set of grid vertices in $\mathcal{T}_{h}$ and $\phi_{h,k}\in V_{h}$ the hat function associated to $a_{h,k}$. Then $\Omega=\cup_{k=1}^{N_{h}}\Omega_{h,k}$ with $\Omega_{h,k}={\rm supp}(\phi_{h,k})$ is an overlapping domain decomposition and $H_{0}^{1}(\Omega)=V_{h}+\sum_{k=1}^{N_{h}}H_{0}^{1}(\Omega_{h,k})$ is a stable subspace decomposition. In this case, the estimator $S_{a}$ in Theorem 2.1 leads to
@@ 3.1 Practical Estimators in H(grad)
Let $\mathcal{T}_{h/2}$ be a uniform refinement of $\mathcal{T}_{h}$ and $V=V_{h/2}$ be the finite element space based on the refined mesh $\mathcal{T}_{h/2}$. For simplicial meshes, $\mathcal{T}_{h/2}$ is constructed by dividing each element in $\mathcal{T}_{h}$ into four congruent children in $\mathbb{R}^{2}$ and by dividing each element in $\mathcal{T}_{h}$ into eight children (cf. [18]) in $\mathbb{R}^{3}$.
@@ 3.1.1 Pointwise Smoother-type Error Estimators
The implicit extrapolation technique [45] provides an efficient way to assemble a $\mathcal{P}_{2}$ smoother matrix $\mathbb{S}_{h,\mathcal{P}_{2}}$ via $\mathbb{S}_{h,\mathcal{P}_{2}}=(4/3)\mathbb{S}_{h/2}-(1/3)\mathbb{S}_{h}$, where the matrices $\mathbb{S}_{h/2}$ and $\mathbb{S}_{h}$ represent the $\mathcal{P}_{1}$ smoothers $S_{h/2}$ and $S_{h}$, respectively.
@@ 3.1.2 Localized Error Estimator and Adaptive Algorithm
To guide the local mesh refinement, each $a_{ii}^{-1}r_{i}^{2}$ in $\langle r_{h/2},S^{a}_{h/2}r_{h/2}\rangle$ is equally subdivided (according to the degree ${\rm deg}(a_{h/2,i})$ of the vertex $a_{h/2,i}$ in $\mathcal{T}_{h/2}$) and allocated to all elements in $\mathcal{T}_{h/2}$ touching the vertex $a_{h/2,i}\in\mathcal{V}_{h/2}$, see Figure 3.1 for example. The local estimator for each $T\in\mathcal{T}_{h}$ reads
$\mathcal{E}_{T}=\left(\sum_{t\subset T,t\in\mathcal{T}_{h/2}}\sum_{a_{h/2,i}\in\bar{t}}\frac{1}{{\rm deg}(a_{h/2,i})}a_{ii}^{-1}r_{ii}^{2}\right)^{\frac{1}{2}}.$}
@@
7: Mark a subset $\mathcal{M}_{h}\subset\mathcal{T}_{h}$ based on Dörfler’s criterion with $\theta\in(0,1)$: $\sum_{T\in\mathcal{M}}\mathcal{E}_{T}^{2}\geq\theta^{2}\sum_{T\in\mathcal{T}_{h}}\mathcal{E}_{T}^{2}$;
@@
In all numerical experiments, the marking parameter $\theta=0.5$ in Algorithm 3.1 will be used. Although the PDE, the domain, the finite element space $V_{h}$, the residual $r_{h}\in V_{h}^{*}$, and the estimator $\mathcal{E}_{T}$ may vary from case to case, we adhere to the general adaptive workflow outlined in Algorithm 3.1. The quantity shown in parentheses in each figure legend is the effectivity ratio, i.e., the value of estimator/error.
@@ 3.1.3 Numerical Experiments in H(grad)
On the L-shaped domain $\Omega=(-1,1)^{2}\backslash([0,1)\times[-1,0))$, we test the performance of the proposed a posteriori error estimates for the FEM (2.9b). The exact solution is $u=r^{2/3}\sin(2\theta/3)$ with singularity near the re-entrant corner of $\Omega$, where $(r,\theta)$ is the polar coordinate near the origin. All quantities are computed on adaptively generated meshes by Algorithm 3.1. The smoother-type a posteriori error estimates as well as computable discrete approximations of implicit estimators (2.10) and (2.11) are compared to the explicit residual error estimator
$\mathcal{E}_{\rm res}=\Big(\sum_{T\in\mathcal{T}_{h}}h_{T}^{2}\|f+\Delta u_{h}\|_{L^{2}(T)}^{2}+\sum_{E\text{ is an interior edge in }\mathcal{T}_{h}}h_{E}\|\llbracket\partial_{n}u_{h}\rrbracket\|^{2}_{L^{2}(E)}\Big)^{\frac{1}{2}},$ (3.7)
@@
Remark 3.2 (Inhomogeneous Dirichlet boundary condition).
For inhomogeneous and piecewise polynomial boundary data $u|_{\partial\Omega}=g_{h}\in V_{h}|_{\partial\Omega}$, the finite element solution $u_{h}\in V_{h}+g_{h}$ exactly matches $u$ and $u_{h/2}$ on the boundary. As a result, $u-u_{h}\in H_{0}^{1}(\Omega)={\rm Dom}(-\Delta)$ and $u_{h/2}-u_{h}\in V_{h/2}={\rm Dom}(A_{h/2})$ are contained in the domain of $\Delta:H_{0}^{1}(\Omega)\rightarrow H^{-1}(\Omega)$ and $A_{h/2}$, respectively. In this case, the same smoothers for $\Delta$ and $A_{h/2}$ in Sections 2.2 and 3.1 are applied to the residual to obtain smoother-type a posteriori error estimates. For general boundary data $u|_{\partial\Omega}=g$, the data oscillation $\big(\sum_{{\rm edge}~E\subset\partial\Omega}h_{E}\|g-g_{h}\|_{L^{2}(E)}^{2}\big)^{\frac{1}{2}}$ is added to the error estimator, where $g_{h}\in V_{h}|_{\partial\Omega}$ is an approximation of $g$.
@@
Smoother-type a posteriori error estimates could be directly generalized to higher order finite elements. Let $S_{\mathcal{P}_{p}}^{a}:V_{\mathcal{P}_{p}}^{*}\rightarrow V_{\mathcal{P}_{p}}$ and $S^{m}_{\mathcal{P}_{p}}:V_{\mathcal{P}_{p}}^{*}\rightarrow V_{\mathcal{P}_{p}}$ be the operator corresponding to the inverse of the diagonal and the inverse of the upper triangular part of the $\mathcal{P}_{p}$ finite element matrix on $\mathcal{T}_{h/2}$, respectively. With $r=r_{h/2}$, the polynomial-degree-robustness ($p$-robustness) of smoother-type estimators $\langle r,S_{\mathcal{P}_{p}}^{a}r\rangle^{\frac{1}{2}}$ and $|S_{\mathcal{P}_{p}}^{a}r|_{H^{1}(\Omega)}$ is numerically demonstrated in Figure 3.3. In addition, Table 3.1 reports effectivity ratios of the residual estimator $\mathcal{E}_{\rm res}$ and the equilibrated estimator $\mathcal{E}_{\rm eq}$ in [36] for controlling the $\mathcal{P}_{p}$ finite element errors with $1\leq p\leq 7$. The numerical results confirm that $\mathcal{E}_{\rm res}$ is not $p$-robust: $\mathcal{E}_{\rm res}/|u-u_{h}|_{H^{1}(\Omega)}$ grows at least linearly in $p$ increases. In contrast, $\mathcal{E}_{\mathrm{eq}}$ provides sharp effectivity ratios close to 1, albeit at the cost of a more involved, implicit patchwise flux reconstruction.
Inspired by multigrid methods that perform multiple smoothing on each level, we consider the smoother-type error estimator $|S^{m,\text{iter}}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$, which corresponds to applying iter steps of Gauss–Seidel smoothing to the residual. Table 3.2 shows that raising iter from 1 to 4 yields only a modest gain in the effectivity ratio. Therefore, we adopt a single smoothing pass in all subsequent experiments.
@@
$\mathcal{E}_{{\rm res},L^{2}}=\Big(\sum_{T\in\mathcal{T}_{h}}h_{T}^{4}\|f+\Delta u_{h}\|_{L^{2}(T)}^{2}+\sum_{E\text{ is an interior edge}}h_{E}^{3}\|\llbracket\partial_{n}u_{h}\rrbracket\|^{2}_{L^{2}(E)}\Big)^{\frac{1}{2}}$
@@
Figure 3.4: Adaptive error estimator convergence in the $L^{2}$ norm for the Poisson equation in fine-coarse layers (left) and $\mathcal{P}_{2}$-$\mathcal{P}_{1}$ layers (right).
Table 3.1: Effectivity ratios of smoother-type $\langle r,S^{a}_{\mathcal{P}_{p}}r\rangle^{1/2}$ and $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$, residual $\mathcal{E}_{\rm{res}}$, and equilibrated $\mathcal{E}_{\mathrm{eq}}$ error estimators.
@@
Table 3.2: Effectivity ratios of the smoother-type error estimator $|S^{m,\text{iter}}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$ based on $1\leq\text{iter}\leq 4$ steps of Gauss–Seidel smoothing operations.

### v3

@@ Abstract
This work develops user-friendly a posteriori error estimates of finite element methods, based on smoothers of linear iterative solvers. The proposed method employs simple smoothers, such as Jacobi or Gauss–Seidel iteration, on an auxiliary finer mesh to process the finite element residual for a posteriori error control. The implementation has linear complexity and requires only a coarse-to-fine prolongation operator. For symmetric problems, we prove the reliability and efficiency of smoother-type error estimators under a saturation assumption. Numerical experiments for various PDEs demonstrate that the proposed smoother-type error estimators outperform residual-type estimators in accuracy and exhibit robustness with respect to physical parameters and polynomial degrees.
@@ 1 Introduction
A posteriori error estimation of Finite Element Methods (FEMs) is also closely related to iterative linear solvers because both methods aim at developing computable approximations to the action of solution operators on some residuals. For instance, the idea of hierarchical basis solvers motivated the development of hierarchical basis error estimators in [9, 6]. Multigrid solvers combined with superconvergence are used for constructing a posteriori error estimates in [12, 13]. In [59], residual error estimators are applied to inaccurate intermediate solutions produced by several smoothing passes to save computational cost in AFEMs. It has been shown in [53, 54] that additive operator preconditioning on infinite-dimensional spaces yields implicit and parameter-robust a posteriori error estimates for $H({\rm grad})$, $H({\rm curl})$ and $H({\rm div})$ problems. Conversely, a posteriori error estimation has been applied to design coarse spaces in multigrid methods and to control iterative algebraic errors in domain decomposition methods, see [72, 45, 15].
@@
We shall follow the research line in [53, 54] to derive a posteriori error estimates based on smoothers of iterative linear solvers. Smoothers are simple basic iterations such as Jacobi or Gauss–Seidel method on a fine mesh and are efficient for damping the high frequency component of the FEM error. We shall show that appropriate smoothers directly yield high-quality a posteriori error estimates of FEMs for various PDEs such as the Poisson, curl-curl, Biot, convection-diffusion and Helmholtz equations. In numerical experiments, smoother-based error estimators exhibit sharper effectivity ratios than residual-type a posteriori error estimates. In contrast to the additive preconditioning-type implicit error estimators derived in [53, 54], which involve incomputable local problems, the proposed smoother-type a posteriori error estimators are fully computable. Moreover, this work introduces a new family of error estimators inspired by contractive iteration methods, which are particularly well-suited for handling non-symmetric problems. For Lagrange finite element discretizations, the smoother-type error estimators are explicit, as they avoid solving patchwise local problems.
@@
The implementation of the smoother-type a posteriori error estimates on a mesh $\mathcal{T}_{h}$ is similar to a two-grid solver. In particular, we shall make use of a simple smoother $S$, e.g., a Jacobi, block Jacobi or Gauss–Seidel iteration on an auxiliary finer mesh $\mathcal{T}_{h/2}$. Then a smoother-type a posteriori error estimate is of the form $\|Sr\|$ or $\sqrt{\langle r,Sr\rangle}$, where $r$ is the finite element residual transferred to $\mathcal{T}_{h/2}$, $\|\bullet\|$ is the norm of interest and $\langle\bullet,\bullet\rangle$ is the pairing between the primal and dual spaces. The reliability and efficiency of our a posteriori error estimates are confirmed under a saturation assumption. Using the same assumption, adaptive finite and boundary element methods driven by the difference $\|u_{h}-u_{h/2}\|$ of finite element solution on nested meshes have been analyzed in [40, 39], while computing the finer mesh solution $u_{h/2}$ is much more expensive than the smoother action $Sr$.
@@
The rest of the paper is organized as follows. In Section 2 we provide preliminaries and set up the abstract framework. In Section 3 we employ smoothers of iterative solvers to derive a posteriori error estimates for positive-definite problems. Section 4 is devoted to the smoother-type a posteriori error estimates of symmetric saddle-point systems. Section 5 illustrates the applications of smoother-type estimators to convection-diffusion and Helmholtz equations. Section 6 is the concluding remarks.
@@ 2.1 Abstract Error Estimators
Throughout this paper, $V_{h}$ is a proper finite element space based on a partition $\mathcal{T}_{h}$ of the underlying domain $\Omega$ for (2.1a). The form $a$ induces a bounded linear operator $A:V\rightarrow V^{*}$ defined as follows:
@@
We adopt the notation $C_{1}\lesssim C_{2}$ provided $C_{1}\leq C_{3}C_{2}$ with $C_{3}$ being a generic constant independent of quantities of interest, e.g., the mesh size parameter $h$ and target functions. We say $C_{1}\eqsim C_{2}$ provided $C_{1}\lesssim C_{2}$, $C_{2}\lesssim C_{1}$. This notation generalizes to comparison between SPD operators in an obvious way. It has been observed in [53] that additive two-level preconditioning yields a reliable and efficient error estimator for SPD problems. For completeness, we include a brief proof.
@@
We remark that $\langle r,S_{a}r\rangle^{\frac{1}{2}}$ is applicable to SPD problems as well as symmetric but indefinite systems posed in a Cartesian product space $V_{1}\times V_{2}\times\cdots$, where $S_{a}={\rm diag}(S_{V_{1}},S_{V_{2}},\ldots)$ is block diagonal with each $S_{V_{i}}$ being an SPD smoother on $V_{i}$, see Section 4. However, $\langle r,S_{a}r\rangle^{\frac{1}{2}}$ could not be applied to strongly non-symmetric problems because $\langle r,S_{a}r\rangle$ is not necessarily nonnegative in such cases. Our new observation is that a contractive iterative solver for $Au=f$
@@
When $A$ is the SPD operator induced by the inner product of $H^{1}(\Omega)$, $H({\rm curl},\Omega)$, $H({\rm div},\Omega)$ or their finite element variants, the iteration (2.7) with $S_{m}$ given by a stable decomposition is uniformly contractive under the norm $\|\bullet\|_{A}$, i.e., $\|I-B_{m}A\|_{A}<1$ is independent of the mesh size parameter $h$.
@@ 2.2 Application to the Poisson Equation
where $V_{h}\subset H_{0}^{1}(\Omega)$ is a finite element space of continuous and piecewise $\mathcal{P}_{1}$ polynomials. Let $\{a_{h,k}\}_{1\leq k\leq N_{h}}$ be the set of grid vertices in $\mathcal{T}_{h}$ and $\phi_{h,k}\in V_{h}$ the hat function associated to $a_{h,k}$. Then $\Omega=\cup_{k=1}^{N_{h}}\Omega_{h,k}$ with $\Omega_{h,k}={\rm supp}(\phi_{h,k})$ is an overlapping domain decomposition and $H_{0}^{1}(\Omega)=V_{h}+\sum_{k=1}^{N_{h}}H_{0}^{1}(\Omega_{h,k})$ is a stable subspace decomposition (cf. [53], Lemma 3.3). In this case, the estimator $S_{a}$ in Theorem 2.1 leads to
@@ 3.1 Practical Estimators in H(grad)
Let $\mathcal{T}_{h/2}$ be a uniform refinement of $\mathcal{T}_{h}$ and $V=V_{h/2}$ be the finite element space based on the refined mesh $\mathcal{T}_{h/2}$. For simplicial meshes, $\mathcal{T}_{h/2}$ is constructed by dividing each element in $\mathcal{T}_{h}$ into four congruent children in $\mathbb{R}^{2}$ or by dividing each element in $\mathcal{T}_{h}$ into eight children (cf. [18]) in $\mathbb{R}^{3}$.
@@ 3.1.1 Pointwise Smoother-type Error Estimators
The implicit extrapolation technique [46] provides an efficient way to assemble a $\mathcal{P}_{2}$ smoother matrix $\mathbb{S}_{h,\mathcal{P}_{2}}$ via $\mathbb{S}_{h,\mathcal{P}_{2}}=(4/3)\mathbb{S}_{h/2}-{(1/3)\tilde{\mathbb{S}}_{h}}$, where $\mathbb{S}_{h/2}$ and $\mathbb{S}_{h}$ represent the $\mathcal{P}_{1}$ smoothers $S_{h/2}$ and $S_{h}$, and $\tilde{\mathbb{S}}_{h}$ is the embedding of $\mathbb{S}_{h}$ into the fine grid matrix size, with zeros assigned to the nodes present in the fine grid $\mathcal{T}_{h/2}$ but not in the coarse grid $\mathcal{T}_{h}$.
@@ 3.1.2 Localized Error Estimator and Adaptive Algorithm
To guide the local mesh refinement, each $a_{ii}^{-1}r_{i}^{2}$ in $\langle r_{h/2},S^{a}_{h/2}r_{h/2}\rangle$ is equally subdivided (according to the vertex degree ${\rm deg}(a_{h/2,i})$, i.e., the number of edges in $\mathcal{T}_{h/2}$ that are incident to $a_{h/2,i}$) and allocated to all elements in $\mathcal{T}_{h/2}$ touching the vertex $a_{h/2,i}\in\mathcal{V}_{h/2}$, see Figure 3.1 for example. The localized estimator for each $T\in\mathcal{T}_{h}$ reads
$\mathcal{E}_{T}=\left(\sum_{t\subset T,t\in\mathcal{T}_{h/2}}\sum_{\begin{subarray}{c}a_{h/2,i}\in\bar{t}\\ a_{h/2,i}\in\mathcal{V}_{h/2}\end{subarray}}\frac{1}{{\rm deg}(a_{h/2,i})}a_{ii}^{-1}r_{i}^{2}\right)^{\frac{1}{2}}.$
@@
7: Mark a $\mathcal{M}_{h}\subset\mathcal{T}_{h}$ with minimal cardinality such that $\sum_{T\in\mathcal{M}_{h}}\mathcal{E}_{T}^{2}\geq\theta^{2}\sum_{T\in\mathcal{T}_{h}}\mathcal{E}_{T}^{2}$,   $\theta\in(0,1)$;
@@
In all numerical experiments, the marking parameter $\theta=0.5$ in Algorithm 3.1 will be used. Although the PDE, the domain, the finite element space $V_{h}$, the residual $r_{h}\in V_{h}^{*}$, and the estimator $\mathcal{E}_{T}$ may vary from case to case, we adhere to the general adaptive workflow outlined in Algorithm 3.1. Numerical experiments were conducted in MATLAB 2024b, with linear systems solved by the direct solver mldivide. The quantity shown in parentheses in each figure legend is the effectivity ratio, i.e., the value of estimator/error on the finest mesh generated by the final AFEM iteration.
@@ 3.1.3 Numerical Experiments in H(grad)
On the L-shaped domain $\Omega=(-1,1)^{2}\backslash([0,1)\times[-1,0))$, we test the performance of the proposed a posteriori error estimates for the FEM (2.9b). The exact solution is $u=r^{2/3}\sin(2\theta/3)$ with singularity near the re-entrant corner of $\Omega$, where $(r,\theta)$ is the polar coordinate near the origin. All quantities are computed on adaptively generated meshes by Algorithm 3.1. The smoother-type a posteriori error estimates as well as computable discrete approximations of implicit estimators (2.10) and (2.11) based on local $\mathcal{P}_{2}$ resolution are compared to the explicit residual error estimator
$\mathcal{E}_{\rm res}=\Big(\sum_{T\in\mathcal{T}_{h}}h_{T}^{2}\|f+\Delta u_{h}\|_{L^{2}(T)}^{2}+\sum_{\text{edge }E\subset\mathring{\Omega}}h_{E}\|\llbracket\partial_{n}u_{h}\rrbracket\|^{2}_{L^{2}(E)}\Big)^{\frac{1}{2}},$ (3.7)
@@
Remark 3.2 (Inhomogeneous essential boundary condition in H(grad)).
For inhomogeneous and piecewise polynomial boundary data $u|_{\partial\Omega}=g_{h}\in V_{h}|_{\partial\Omega}$, the finite element solution $u_{h}\in V_{h}+g_{h}$ exactly matches $u$ and $u_{h/2}$ on the boundary. As a result, $u-u_{h}\in H_{0}^{1}(\Omega)={\rm Dom}(-\Delta)$ and $u_{h/2}-u_{h}\in V_{h/2}={\rm Dom}(A_{h/2})$ are contained in the domain of $\Delta:H_{0}^{1}(\Omega)\rightarrow H^{-1}(\Omega)$ and $A_{h/2}$, respectively. In this case, the same smoothers for $\Delta$ and $A_{h/2}$ in Sections 2.2 and 3.1 are applied to the residual to obtain smoother-type a posteriori error estimates. For general boundary data $u|_{\partial\Omega}=g$, we implement the discrete boundary condition $u_{h}|_{\partial\Omega}=g_{h}$ with $g_{h}\in V_{h}|_{\partial\Omega}$ being the nodal interpolant of $g$. Rigorously speaking, the data oscillation ${\rm osc}_{\partial}=\big(\sum_{{\rm edge}~E\subset\partial\Omega}h_{E}|g-g_{h}|_{H^{1}(E)}^{2}\big)^{\frac{1}{2}}$ for 2D problems should be incorporated into both smoother-type and other classical error estimators to obtain rigorous a posteriori error upper bounds (cf. [58, 38]). In practice, the term $\rm osc$$_{\partial}$ is of higher order than the FEM error and therefore does not significantly influence the performance of AFEMs. For simplicity of presentation, we omit $\rm osc$$_{\partial}$ from all error estimators under comparison in the subsequent experiments.
@@
Smoother-type a posteriori error estimates could be directly generalized to higher order finite elements. Let $S_{\mathcal{P}_{p}}^{a}:V_{\mathcal{P}_{p}}^{*}\rightarrow V_{\mathcal{P}_{p}}$ and $S^{m}_{\mathcal{P}_{p}}:V_{\mathcal{P}_{p}}^{*}\rightarrow V_{\mathcal{P}_{p}}$ be the operators corresponding to the inverse of the diagonal and the inverse of the upper triangular part of the $\mathcal{P}_{p}$ finite element matrix on $\mathcal{T}_{h/2}$, respectively. With $r=r_{h/2}$, the polynomial-degree-robustness ($p$-robustness) of smoother-type estimators $\langle r,S_{\mathcal{P}_{p}}^{a}r\rangle^{\frac{1}{2}}$ and $|S_{\mathcal{P}_{p}}^{a}r|_{H^{1}(\Omega)}$ is numerically demonstrated in Figure 3.3 and Table 3.1. In addition, Table 3.1 reports effectivity ratios of the residual estimator $\mathcal{E}_{\rm res}$ and the equilibrated estimator $\mathcal{E}_{\rm eq}$ in [36] for controlling the $\mathcal{P}_{p}$ finite element errors with $1\leq p\leq 7$. The numerical results confirm that $\mathcal{E}_{\rm res}$ is not $p$-robust: $\mathcal{E}_{\rm res}/|u-u_{h}|_{H^{1}(\Omega)}$ grows at least linearly as $p$ increases. In contrast, $\mathcal{E}_{\mathrm{eq}}$ provides sharp effectivity ratios $\mathcal{E}_{\rm eq}/|u-u_{h}|_{H^{1}(\Omega)}$ close to 1, albeit at the cost of a more involved, implicit patchwise flux reconstruction.
Inspired by multigrid methods that perform multiple smoothing on each level, we consider a smoother-type error estimator $|S^{m,\text{iter}}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$, which corresponds to applying iter steps of Gauss–Seidel smoothing to the residual. Table 3.2 shows that raising iter from 1 to 4 yields only a modest gain in the effectivity ratio. Therefore, we adopt a single smoothing pass in all subsequent experiments.
@@
$\mathcal{E}_{{\rm res},L^{2}}=\Big(\sum_{T\in\mathcal{T}_{h}}h_{T}^{4}\|f+\Delta u_{h}\|_{L^{2}(T)}^{2}+\sum_{\text{edge }E\subset\mathring{\Omega}}h_{E}^{3}\|\llbracket\partial_{n}u_{h}\rrbracket\|^{2}_{L^{2}(E)}\Big)^{\frac{1}{2}}$
@@
Figure 3.4: Adaptive error estimator convergence in the $L^{2}$ norm for the Poisson equation in fine-coarse layers (left) and $\mathcal{P}_{2}$-$\mathcal{P}_{1}$ layers (right), where $N$ is the number of triangles.
Table 3.1: Effectivity ratios of smoother-type $\langle r,S^{a}_{\mathcal{P}_{p}}r\rangle^{1/2}$, $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$ and $|S^{m}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$, residual $\mathcal{E}_{\rm{res}}$, and equilibrated $\mathcal{E}_{\mathrm{eq}}$ error estimators (converged values computed on the finest asymptotic-regime mesh).
@@
Table 3.2: Effectivity ratios of the smoother-type error estimator $|S^{m,\text{iter}}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$ based on $1\leq\text{iter}\leq 4$ steps of Gauss–Seidel smoothing operations (converged values computed on the finest asymptotic-regime mesh).